# HR-INR: Continuous Space-Time Video Super-Resolution via Event Camera

### v1

@@ 4.1 Comparison Experiments
Space-time Super-resolution: We conduct space-time super-resolution comparison experiments on the Adobe240 and GoPro datasets. We categorized the comparison methods into three groups. (I) Frame-based cascade methods: VFI methods, e.g., Super SloMo [22] and DAIN [3] followed by VSR methods, e.g., EDVR [66] and BasicVSR [5]. (II) Fixed STVSR methods: e.g., Zooming Slow-Mo [69] and TMNet [70]. (III) Frame-based C-STVSR methods: VideoINR [10] and MoTIF [8]. The numerical comparison is presented in Tab. 1. It is evident that C-STVSR methods consistently outperform cascade and fixed STVSR methods.Our method achieves the highest performance in both datasets with the smallest model size. In the GoPro dataset, our method improves the center frame by 0.93 $dB$ and 0.0415 SSIM, and on average by 2.09 $dB$ and 0.0598 SSIM compared to the best method, MoTIF [8]. Similarly, in the Adobe240 dataset, our method outperforms MoTIF [8] by 0.63 $dB$ and 0.0407 SSIM for the -center, and on -average by 1.29 $dB$ and 0.0466 SSIM. It is worth noting that the difference between our method and other methods is more pronounced for the -average frames than the -center frame. This observation suggests that our method demonstrates enhanced temporal stability and superior adaptability across varying interpolation intervals, an advantage not shared by previous methods [10, 8]. For more analysis, please refer to Sec. 4.2 and Fig. 8 in appendix.
Table 3: Temporal super-resolution results, i.e.,, VFI, on BS-ERGB dataset [61].
@@
FLAVR [26] - 25.95 - 0.086 20.90 - 0.151
DAIN [3] 24.0 25.20 - 0.067 21.40 - 0.113
Super SloMo [22] 19.8 - - - 22.48 - 0.115
QVI [71] 29.2 - - - 23.20 - 0.110
TimeLens [62] 72.2 28.36 - 0.026 27.58 - 0.031
TimeLens++ [61] 53.9 28.56 - 0.022 27.63 - 0.026
CBMNet [28] 15.4 29.32 0.815 - 28.46 0.806 -
CBMNet-Large [28] 22.2 29.43 0.816 - 28.59 0.808 -
@@
Tab. 2 presents additional comparative experiments for arbitrary spatial and temporal super-resolution. Our method consistently outperforms other methods, even in extreme space-time upsample scales, such as 16×16$\times$16 × temporal upscale and 6×6$\times$6 × spatial upscale. The visualization results, Fig. 1 (c) and Fig. 5, also demonstrate that our method effectively models regional nonlinear motion, e.g., the wheel rotation — a capability not achieved by previous method VideoINR [10]. For more visualization results, please refer to the Suppl. Mat.
Figure 6: 3-skip frame interpolation visualization results in BS-ERGB dataset [61]. Our method accurately captures local non-linear motion (e.g., balls in (1.d) and (2.d)), surpassing TimeLens [61], which exhibits ghosting and holes (green circles). A yellow arrow shows TimeLens’s inaccurate ball positioning.
Separate Comparison of Event-based VFI and VSR: We also compared our method with previous approaches in separate VFI [61, 62, 28] and VSR [38, 23] tasks. The results can be seen in Tab. 3 and Tab. 4 respectively. In the VFI task, our method surpasses TimeLens++ [61] by 1.1 $dB$ for 1-skip and 0.17 $dB$ for 3-skip VFI, and CBMNet [28] by 0.23 $dB$ for 1-skip. Additionally, our model’s size is merely $1/7$ and $1/3$ that of TimeLens++ [61] and CBMNet [28], respectively. In the visualization aspect, our approach excels in modeling non-linear motion and long-term dependencies, as evident in Fig. 6 in appendix. It precisely forecasts the positions of small balls (Fig. 6 (1.d)) and the soccer ball (Fig. 6 (2.d)) at intermediate timestamps, outperforming previous methods.
For the VSR task, our method outperforms EG-VSR [38] by 1.03 $dB$ and 0.447 SSIM for 4 $\times$ super-resolution, and by 3.02 $dB$ and 0.0134 SSIM for 2 $\times$ super-resolution. Fig. 7 demonstrates our model’s effective VSR performance on the real-dataset CED, highlighting its ability to reproduce sharp edges and robustness to noise. For more visualization results, see the Suppl. Mat.
Table 4: Spatial super-resolution results on CED dataset. * denotes the values obtained from the official pre-trained models.
      4×4$\times$4 ×       2×2$\times$2 ×
@@
      DUF [24]       1.90       ✗       24.43       0.8177       31.83       0.9183
      TDAN [60]       1.97       ✗       27.88       0.8231       33.74       0.9398
      SOF [64]       1.00       ✗       27.00       0.8050       31.84       0.9226
      RBPN [19]       12.18       ✗       29.80       0.8975       36.66       0.9754
      BasicVSR++ [6]       7.30       ✗       14.76       0.1641       -       -
      VideoINR* [10]       11.31       ✗       25.53       0.7871       26.77       0.7938
      E-VSR [23]       412.42       ✓       30.15       0.9052       37.32       0.9783
      EG-VSR [38]       2.45       ✓       31.12       0.9211       38.69       0.9771
      HR-INR (Our)       8.27       ✓       32.15       0.9658       42.01       0.9905
Figure 7: 4×4$\times$4 × video super-resolution visualization results in CED dataset [52].
Table 5: Ablation studies in Adobe-Average [56] (4×4$\times$4 × and 7-skip). The †  symbol marks the line for comparison with other lines.
Case Events TPR Temporal Embedding Temporal Dim ($C_{t}$) Input Frames PSNR $\uparrow$ SSIM $\uparrow$
Case#1 Learning 640 4 26.84 ( $-4.27$) 0.8366 ( $-0.0850$)
Case#2 Learning 640 4 29.69 ( $-1.41$) 0.8974 ( $-0.0242$)
@@
Case#4 Sinusoid 640 4 30.42 ( $-0.69$) 0.9120 ( $-0.0096$)
Case#5 Learning 320 4 28.44 ( $-2.67$) 0.8700 ( $-0.0516$)
Case#6 Learning 640 2 30.41 ( $-0.70$) 0.9151 ( $-0.0065$)
Case#7 Learning 640 3 30.72 ( $-0.39$) 0.9174 ( $-0.0042$)-
Table 6: Ablation studies for TPR levels and moments in Adobe-Average [56] (4×4$\times$4 × and 7-skip). The †  symbol marks the line for comparison with other lines.
      Case       TPR Level ($L$)       TPR Moments ($M_{p}$)       PSNR $\uparrow$       SSIM $\uparrow$
      Case#1       3       3       29.93 ( $-1.18$)       0.9011 ( $-0.0205$)
      Case#2       5       3       30.32 ( $-0.79$)       0.9165 ( $-0.0051$)
      Case#3       7       3       30.78 ( $-0.33$)       0.9187 ( $-0.0029$)
      Case#4 †       7       9       31.11       0.9216
      Case#5       7       18       31.18 ( $+0.07$)       0.9228 ( $+0.0012$)
Table 7: Ablation in BS-ERGB (TimeLens++) dataset [61].
@@
29.66 ( $+1.41$) 0.8281 ( $+0.0094$) 0.011 ( $-0.007$) 28.59 ( $+1.94$) 0.8140 ( $+0.0273$) 0.021 ( $-0.018$)
4.2 Ablation and Analytical Studies
Ablation and analytical studies conducted on the Adobe240 [56] and BS-ERGB [61] datasets unveiled several critical insights. On the Adobe240 dataset, we executed simultaneous 7-skip VFI and 4×4$\times$4 × VSR tests, as shown in Tab. 5 and Tab. 6, while on the BS-ERGB dataset, 1-skip and 3-skip VFI were performed in Tab. 7. Events Gain: Tab. 5-Case#1 shows that with events input replaced as zero and unchanged network architecture, PSNR and SSIM significantly drop, highlighting the importance of events for temporal motion learning. Adding events alone substantially raised PSNR by $2.85dB$ and SSIM by $0.06$. Event TPR Enhancement: Incorporating event TPR enhances performance, with PSNR and SSIM increasing with layer count, peaking at seven layers, as shown in Tab. 5-Case#2 and Tab. 6. However, expanding ETPR moments from three to seven only improves PSNR by $0.33dB$, with minimal gains beyond that, suggesting a plateau in performance benefits. The TPR also demonstrates enhancement on the BS-ERGB dataset, as shown in Tab. 7, yielding an increase of $0.92dB$ for 1-skip and an improvement of $1.13dB$ for 3-skip. This indicates that the benefits of TPR become more pronounced with the increase in the number of skips. Time Embedding Method: Table 5-Case#3#4 shows Sinusoid embedding’s results. It’s outperformed by learning-based methods, confirmed by previous research [59, 50, 82, 1], due to their superior high-frequency information capture. Temporal Dimension Impact: The INR temporal dimension significantly influences performance. Lowering the dimensions from 640 to 320 degrades performance in Tab. 5-Case#3#5, suggesting a reduction in temporal detail capture. Conversely, expanding the dimension to 960 poses instability risks (e.g., $NAN$ errors). This highlights the need to balance dimensionality and training stability. Input Frames: Tab. 5-Case#3#6#7 illustrates the impact of varying input frame counts on the final results. We observed a performance decrease of 0.70 $dB$ and 0.39 $dB$ when inputting two and three frames, respectively, compared to four frames. This indicates a clear advantage of multi-frame inputs in modeling longer-term dependencies. Moreover, even with two frames, our method also outperforms previous works [10, 8].

### v2

@@ 4.2 Comparison Experiments
Since our method is the first event-based C-STVSR approach, we compare it with both frame-based methods and cascade methods for C-STVSR, as discussed in Sec. 4.2.1. Additionally, our method also demonstrates superior performance in individual tasks, VFI and VSR. The comparisons for these individual tasks are discussed in Sec. 4.2.2.
Figure 9: Visualization results on a real-world dataset in a large object long-distance motion scene. (a) and (b) represent the input frames, (c) shows the corresponding events, (d) is the result of the TimeLens method (TL), (e) shows the interpolation by CBMNet (CL), (f) presents the result of our method, and (j) is the ground truth (GT), and (h) displays the local events at each time step.
4.2.1 Space-time Super-resolution
We conduct space-time super-resolution comparison experiments on the Adobe240 and GoPro datasets. We categorized the comparison methods into three groups. (I) Cascade methods: VFI methods, e.g., Super SloMo (Jiang et al, 2018), DAIN (Bao et al, 2019a) and TimeLens (Tulyakov et al, 2021) followed by VSR methods, e.g., EDVR (Wang et al, 2019), BasicVSR (Chan et al, 2021) and EG-VSR (Lu et al, 2023). (II) Fixed STVSR methods: e.g., Zooming Slow-Mo (Xiang et al, 2020) and TMNet (Xu et al, 2021). (III) Frame-based C-STVSR methods: VideoINR (Chen et al, 2022) and MoTIF (Chen et al, 2023b).
The numerical comparison is presented in Tab. 1. It is evident that C-STVSR methods consistently outperform cascade and fixed STVSR methods. Our method achieves the highest performance in both datasets with the smallest model size. In the GoPro dataset, our method improves the center frame by 0.93 $dB$ and 0.0415 SSIM, and on average by 2.09 $dB$ and 0.0598 SSIM compared to the best method, MoTIF (Chen et al, 2023b). Similarly, in the Adobe240 dataset, our method outperforms MoTIF (Chen et al, 2023b) by 0.63 $dB$ and 0.0407 SSIM for the -center, and on -average by 1.29 $dB$ and 0.0466 SSIM. It is worth noting that the difference between our method and other methods is more pronounced for the -average frames than the -center frame. This observation suggests that our method demonstrates enhanced temporal stability, as shown in Fig. 7, and superior adaptability across varying interpolation intervals, an advantage not shared by previous methods (Chen et al, 2022, 2023b).
Tab. 2 presents additional comparative experiments for arbitrary spatial and temporal super-resolution. Our method consistently outperforms other methods, even in extreme space-time upsample scales, such as $16\times$ temporal upscale and $6\times$ spatial upscale. The visualization results, Fig. 1 (c) and Fig. 6, also demonstrate that our method effectively models regional nonlinear motion, e.g., the wheel rotation — a capability not achieved by previous method VideoINR (Chen et al, 2022). For more video visualization results, please refer to the Supplementary Materials.
Stability Across Temporal Sequences: Our method demonstrates temporal stability, accurately estimating motion states at each time point during frame interpolation, as shown in Fig. 7. This is the key reason why our method outperforms previous approaches. Specifically, whether in the GoPro-Average and Adobe-Average of Tab. 1, or the 12-skip and 16-skip tests of Tab. 2, our method significantly outperforms previous methods by at least $1.2dB$ with $4\times$ VSR.
Figure 10: Visualization results on a real-world dataset in a small object long-distance motion scene. (a) and (b) represent the input frames, (c) shows the corresponding events, (d) is the result of the TimeLens method (TL), (e) shows the interpolation by CBMNet-Large (CL), (f) presents the result of our method, and (j) is the ground truth (GT), and (h) displays the local events at each time step.
Table 3: Temporal super-resolution results, i.e.,, VFI, on BS-ERGB dataset (Tulyakov et al, 2022).
@@ 4.2.1 Space-time Super-resolution
FLAVR (Kalluri et al, 2023) - 25.95 - 0.086 20.90 - 0.151
DAIN (Bao et al, 2019a) 24.0 25.20 - 0.067 21.40 - 0.113
Super SloMo (Jiang et al, 2018) 19.8 - - - 22.48 - 0.115
QVI (Xu et al, 2019) 29.2 - - - 23.20 - 0.110
TimeLens (Tulyakov et al, 2021) 72.2 28.36 - 0.026 27.58 - 0.031
TimeLens++ (Tulyakov et al, 2022) 53.9 28.56 - 0.022 27.63 - 0.026
CBMNet (Kim et al, 2023) 15.4 29.32 0.815 - 28.46 0.806 -
CBMNet-Large (Kim et al, 2023) 22.2 29.43 0.816 - 28.59 0.808 -
@@
Fig. 7 shows the relationship between timestamps and PSNR, validating the greater stability of our method compared to VideoINR  (Chen et al, 2022), especially around the $0.5s$ mark, where VideoINR experiences a notable decrease. This observation is also reflected in the visualization results of Fig. 16 and the following Fig. 8910. Whether using real-world or simulated data, our method demonstrates superior temporal consistency, exhibiting fewer artifacts compared to previous methods such as VideoINR (Chen et al, 2022), TimeLens (Tulyakov et al, 2021) and CMB-Net (Kim et al, 2023). This also highlights the real-world effectiveness of our method’s temporal stability. The main reason for this is that previous methods (Chen et al, 2022, 2023b; Tulyakov et al, 2021) suffer from instability in motion estimation at timestamps far from the reference frames (at 0 and 11 timestamp), which negatively impacts shape and edge estimation for frame interpolation. In contrast, our method benefits from the intermediate motion details provided by events TPR and the model’s ability to capture long temporal sequences, enabling it to maintain higher stability across the temporal sequence and deliver superior performance.
Figure 11: $4\times$ video super-resolution visualization results in CED dataset (Scheerlinck et al, 2019b).
Table 4: Spatial super-resolution results on CED (Scheerlinck et al, 2019b). * denotes values from pre-trained models.
$4\times$ $2\times$
@@
TDAN (Tian et al, 2020) 1.97 27.88 0.8231 33.74 0.9398
SOF (Wang et al, 2020) 1.00 27.00 0.8050 31.84 0.9226
RBPN (Haris et al, 2019) 12.18 29.80 0.8975 36.66 0.9754
VideoINR* (Chen et al, 2022) 11.31 25.53 0.7871 26.77 0.7938
E-VSR (Jing et al, 2021) 412.42 30.15 0.9052 37.32 0.9783
EG-VSR (Lu et al, 2023) 2.45 31.12 0.9211 38.69 0.9771
HR-INR (Our) 8.27 32.15 0.9658 42.01 0.9905
4.2.2 Separate Comparison of Event-based VFI and VSR
We also compared our method with previous approaches in separate VFI (Tulyakov et al, 2022, 2021; Kim et al, 2023) and VSR (Lu et al, 2023; Jing et al, 2021) tasks. The results can be seen in Tab. 3 and Tab. 4 respectively.
Comparison with Event-based VFI: In the VFI task, our method surpasses TimeLens++ (Tulyakov et al, 2022) by 1.1 $dB$ for 1-skip and 0.17 $dB$ for 3-skip VFI, and CBMNet (Kim et al, 2023) by 0.23 $dB$ for 1-skip in Tab. 3. Additionally, our model’s size is merely $1/7$ and $1/3$ that of TimeLens++ (Tulyakov et al, 2022) and CBMNet (Kim et al, 2023), respectively. In the visualization aspect, our approach excels in modeling non-linear motion and long-term dependencies, as evident in Fig. 8, where the subject is juggling, the motion of small, fast-moving objects (balls) is challenging to capture. Our method precisely forecasts the positions of small balls (Fig. 8 (1.d)) and the soccer ball (Fig. 8 (2.d)) at intermediate timestamps. Competing methods exhibit noticeable ghosting and fail to preserve the spatial consistency of the balls. This demonstrates our model’s ability to handle high-speed motion effectively by leveraging its dual-branch feature extraction.
Table 5: Ablation studies in Adobe-Average (Su et al, 2017) ($4\times$ and 7-skip). The †  symbol marks the line for comparison with other lines.
Case Events TPR
Temporal
Embedding
Temporal
Dim ($C_{t}$)
Input
Frames
PSNR $\uparrow$
SSIM $\uparrow$
Case#1 Learning 640 4 26.84 ( 4.27-4.27) 0.8366 ( 0.0850-0.0850)
Case#2 Learning 640 4 29.69 ( 1.41-1.41) 0.8974 ( 0.0242-0.0242)
@@ 4.2.2 Separate Comparison of Event-based VFI and VSR
Case#4 Sinusoid 640 4 30.42 ( 0.69-0.69) 0.9120 ( 0.0096-0.0096)
Case#5 Learning 320 4 28.44 ( 2.67-2.67) 0.8700 ( 0.0516-0.0516)
Case#6 Learning 640 2 30.41 ( 0.70-0.70) 0.9151 ( 0.0065-0.0065)
Case#7 Learning 640 3 30.72 ( 0.39-0.39) 0.9174 ( 0.0042-0.0042)-
Furthermore, we also explore the performance of our method in scenarios involving motions of objects of different sizes. Fig. 9, 10 demonstrate the superiority of our method in recovering large-scale and long-distance motions across various real-world scenarios. These results highlight the effectiveness of our approach in fusing both regional and holistic information to handle complex motion dynamics, while simultaneously reducing ghosting artifacts that are commonly seen in competing methods. Fig. 9 shows large object long-distance motion, such as the horse and rider jumping, our method shows clear advantages. The silhouette of the horse and the rider’s position are reconstructed with remarkable clarity and consistency. In contrast, competing methods introduce significant blurring and fail to preserve the integrity of the subject’s shape, revealing their limitations in handling large-scale motion over time. Fig. 10 shows small object long-distance motion. Our method excels in reconstructing fine-grained motion details, such as the movement of the horse’s legs. Unlike CMB-Net and TimeLens, which either fail to capture these intricate details or introduce substantial motion blur, our approach accurately restores the distinct positions of the horse’s hooves. This ability arises from the integration of local motion features captured by the regional branch, along with the long-term temporal dependencies modeled by the holistic branch. Further analysis can be found in Sec. 4.3.
Across all scenarios, our method generates outputs with fewer ghosting artifacts. The combination of regional and holistic features allows our approach to capture both short-term and long-term motion dependencies, resulting in more realistic and temporally coherent video outputs. These results validate the robustness and generalizability of our method across diverse motion patterns and scales.
Comparison with Event-based VSR: For the VSR task, our method outperforms EG-VSR (Lu et al, 2023) by 1.03 $dB$ and 0.447 SSIM for 4 $\times$ super-resolution, and by 3.02 $dB$ and 0.0134 SSIM for 2 $\times$ super-resolution. As shown in Fig. 10, our method produces visually sharper results compared to EG-VSR and other baseline methods. In Fig.  11, we present video super-resolution visualization results on the CED dataset (Scheerlinck et al, 2019b), where our approach effectively restores fine details, such as the text on the building and edges around structures, which are blurred or distorted in competing methods. The corresponding error map (Fig.  11 (g)) further demonstrates our method’s reduced errors, especially around high-contrast regions.
Figure 12: Feature visualization on real data (Tulyakov et al, 2022): (f) shows the holistic feature, $F^{g}$, derived from multiple frames and events; (g) depict the regional features ($F_{t}^{g}$), highlighting the capability to capture local motion.
4.3 Ablation and Analytical Studies
Ablation and analytical studies conducted on the Adobe240 (Su et al, 2017) and BS-ERGB (Tulyakov et al, 2022) datasets unveiled several critical insights. On the Adobe240 dataset, we executed simultaneous 7-skip VFI and $4\times$ VSR tests, as shown in Tab. 5 and Tab. 6, while on the BS-ERGB dataset, 1-skip and 3-skip VFI were performed in Tab. 7.
Table 6: Ablation studies for TPR levels and moments in Adobe-Average Su et al (2017) ($4\times$ and 7-skip). The †  symbol marks the line for comparison with other lines. ”Captured Moment” refers to the temporal resolution of the last layer of the TPR, which is calculated by Eq. 4.
Case TPR Level ($L$) TPR Moments ($M_{p}$) Captured Moment PSNR $\uparrow$ SSIM $\uparrow$
Case#1 3 3 1 / 81 29.93 ( 1.18-1.18) 0.9011 ( 0.0205-0.0205)
Case#2 5 3 1 / 729 30.32 ( 0.79-0.79) 0.9165 ( 0.0051-0.0051)
Case#3 7 3 1 / 6561 30.78 ( 0.33-0.33) 0.9187 ( 0.0029-0.0029)
Case#4 † 7 9 1 / 19683 31.11 0.9216
Case#5 7 18 1 / 39366 31.18 ( +0.07+0.07) 0.9228 ( +0.0012+0.0012)
Table 7: Ablation in BS-ERGB (TimeLens++) dataset Tulyakov et al (2022).
@@ 4.3 Ablation and Analytical Studies
29.66 ( +1.41+1.41) 0.8281 ( +0.0094+0.0094) 0.011 ( 0.007-0.007) 28.59 ( +1.94+1.94) 0.8140 ( +0.0273+0.0273) 0.021 ( 0.018-0.018)
Events Gain: Tab. 5-Case#1 shows that with events input replaced as zero and unchanged network architecture, PSNR and SSIM significantly drop, highlighting the importance of events for temporal motion learning. Adding events alone substantially raised PSNR by $2.85dB$ and SSIM by $0.06$.
Event TPR Analysis: Incorporating event TPR significantly enhances performance, with both PSNR and SSIM improving as the layer count increases, peaking at seven layers, as shown in Tab. 5 and Tab. 6. Specifically, in Tab. 6, as the TRP Level $L$ increases, the moments captured by the TPR become more precise. For example, when the TRP Level rises from 3 in Tab. 6-Case#1, to 7 in Tab. 6-Case#4, the PSNR shows an increase of approximately 1.18. Moreover, the model’s performance further improves with the increase in TPR Moments $M_{p}$. However, it is important to note that when both $L$ and $M_{p}$ are relatively high, the performance improvement tends to plateau. Additionally, the TPR enhancement is also evident on the BS-ERGB dataset, as shown in Tab. 7, where it yields an increase of $0.92$ dB for 1-skip and an improvement of $1.13$ dB for 3-skip.
Analysis of Event TPR in Visualization: Building upon the performance analysis of TPR, we now examine its effectiveness in visualization. As shown in Fig. 12, TPR’s regional features emphasize localized motion, demonstrating its ability to capture fine-grained, short-term dynamics. This capability highlights TPR as a novel event representation for millisecond-scale motion at high temporal precision. For instance, the regional features (Fig.12(h)) focus on localized, short-term motion, while the holistic features (Fig.12(d)) capture the broader, long-term context. Together, these two types of features allow our method to effectively separate dynamic motion from static background information. In the Global Motion Scene, as shown in Fig. 12, where vehicles move against a stationary urban background, the regional features capture the intricate, millisecond-level motion of the vehicles, while the holistic features retain the overall structure of the scene, including static elements such as traffic signs and buildings. This dual representation enables our method to handle complex dynamic scenes seamlessly, where fast-moving objects coexist with stationary elements. In the Local Motion Scene, as shown in Fig. 12, the visualization of a basketball player dribbling the ball reveals that the holistic features primarily encode static background details like trees and stationary objects, while the regional features focus on the fast, dynamic motion of the basketball. This demonstrates TPR’s strength in isolating short-term motion, even for high-speed moving objects. These visualizations provide strong evidence of TPR’s capability to capture fine-grained, short-term motion through its hierarchical structure. By focusing on temporal granularity, event data complements frame-based information and enables the extraction of rich local motion features. We also encourage readers to refer to the supplementary videos for further demonstrations, where these cases are presented in more detail, showcasing how TPR integrates both event and frame data to effectively handle complex motion patterns. This visualization underscores the unique advantages of TPR, combining regional and holistic information to enable robust spatiotemporal super-resolution.
Time Embedding Method: Tab. 5-Case#3#4 presents the results for sinusoid embeddings. The learning-based methods perform better due to their superior ability to capture high-frequency information (Ramasinghe and Lucey, 2023; Attal et al, 2022). This advantage is further demonstrated in Fig. 13, which visualizes the time embedding features. The learning-based embeddings not only effectively capture periodic positional representations but also offer a more comprehensive expression of exposure information.
Temporal Dimension Impact: The INR temporal dimension significantly influences performance. Lowering the dimensions from 640 to 320 degrades performance in Tab. 5-Case#3#5, suggesting a reduction in temporal detail capture. Conversely, expanding the dimension to 960 poses instability risks (e.g., $NAN$ errors). This highlights the need to balance dimensionality and training stability.
Figure 14: Visualization of Holistic Features for different frames as input.
Analysis of Input Frames and Long-distance Modeling: Tab. 5-Case#3#6#7 illustrates the impact of varying input frame counts on the final results. We observed a performance decrease of 0.70 $dB$ and 0.39 $dB$ when inputting two and three frames, respectively, compared to four frames. This indicates a clear advantage of multi-frame inputs in modeling longer-term dependencies. Moreover, even with two frames, our method also outperforms previous works (Chen et al, 2022, 2023b). Fig. 14 compares the outputs of two models: one using two input frames (0, 1) and another using four input frames (-1, 0, 1, 2). With two input frames, the model struggles to capture complex motion, as seen in the less accurate predictions of the basketball’s trajectory in Fig.14 (b). In contrast, the four-frame model leverages additional temporal information to model longer-term motion, producing smoother and more accurate results, as shown in Fig.14 (c). The holistic feature visualizations further highlight the richer temporal dependencies captured by the four-frame model, allowing it to handle challenging motion scenarios more effectively. This analysis underscores the strength of our method in utilizing multi-frame inputs to improve temporal coherence and enhance interpolation quality, particularly for long-distance or rapid motion.
Figure 15: Comparison of total and average time for $34\times$ frame interpolation by different methods. Our method takes less time than TimeLens (Tulyakov et al, 2021), but slightly more time than VideoINR (Chen et al, 2022).
Inference time analysis In Fig. 13, we analyze the inference times of three different methods. Both our method and VideoINR achieve an average frame time of less than 100 $ms$ for $34\times$ frame interpolation. In contrast, TimeLens (Tulyakov et al, 2021) has an average frame time of 187 $ms$, which is more than double that of our method. The tests were conducted on a high-performance computer, and each method was tested 30 times, with the final inference time being the average of these 30 trials.
Bad case analysis We have observed that our method has certain limitations in some cases (Fig. 8). For example, when restoring color information, although our model can accurately reconstruct the contours of objects, the color information is often distorted or missing. This issue primarily arises due to the lack of color information in the event stream. We believe that with future advancements in color event technology, this problem will be effectively addressed.